# Stability-Informed Initialization of Neural Ordinary Differential Equations

### v1

@@ 6.1 Pixel-Level Image Classification
Drawing inspiration from (Le et al., 2015; Gu et al., 2022), the original MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets are modified for a sequential classification task. Rather than presenting the entire image at once, the adaptation introduces the images to the model pixel by pixel sequentially.
@@ Results
For each dataset, $10$ model instances for each initializing technique were trained. The resulting accuracy is summarized in Table 1, reporting the mean and standard deviation. Here, pMNIST refers to the permuted MNIST variant (Goodfellow et al., 2013). It is clear from the results that using the proposed initializing technique provides positive improvements in test accuracy. A notable impact of employing SII is the clear enhancement in training efficiency, achieving convergence at greater speed. This effect is depicted in Figure 6, where the test acccuracy curves are illustrated across the training epochs.
@@ 6.3 Multivariate Time-Series Forecasting
where $\hat{\bm{y}}\in\mathbb{R}^{{d_{y}}}$ is the measurement prediction, and $\bm{u}\in\mathbb{R}^{{d_{u}}}$ is a vector of input measurement signals. Importantly, the system is partly described by a state-transition function (19a), with unknown states and dynamics. In order to compute predictions according to (19b), the latent states $\bm{z}\in\mathbb{R}^{{d_{z}}}$ must be solved for, which requires that $\bm{z}_{0}$ is known. Inspired by Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014), we employ an encoder network $g_{\phi}$ that outputs the variational posterior over the initial state $\bm{z}_{0}\sim q_{\phi}(\bm{z}_{0}|\bm{u}_{0})$ and parametrize it as diagonal Gaussian with learnable parameters.
@@ Results
A total of $3\times 10$ different models per initialization method were trained, consisting of $3$ random data splits (to mitigate some selection bias (Cawley & Talbot, 2010)) with $10$ different model configurations each. Test performance is presented in Table 3. Using SII results in an approximately $50$% improvement in test performance compared to the default initialization. This enhancement is further supported by the faster convergence rate, as depicted in Figure 7.
@@ 8 Conclusions
This paper analyses how the stability properties of the system and stability regions of the chosen numerical solver affect the training and prediction performance of neural ODEs. Further, it is illustrated how standard techniques for the initialization of the network, without considering the stability properties of the system and solver, may lead to slow training and suboptimal performance. Based on this, a general stability-informed initialization (SII) method is developed that adapts to the stability region of a chosen solver. The effectiveness of the approach is demonstrated in various machine-learning tasks, including successful applications to real-world measurement data. In all cases, increased efficiency of training and model performance are demonstrated.
@@ Acknowledgements
This research was supported by the Strategic Research Area at Linköping-Lund in Information Technology (ELLIIT) and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre.

### v3

@@ 6.1 Pixel-Level Image Classification
Drawing inspiration from (Le et al., 2015; Gu et al., 2022), the original MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets are modified for a sequential classification task. This modification involves presenting images to the model one pixel at a time, requiring the model to classify the image only after all pixels have been observed. These tasks have been popularized as benchmarks for long-range dependency problems.
@@ Results
For each dataset, $10$ model instances for each initializing technique were trained. The resulting accuracy is summarized in Table 1, reporting the mean and standard deviation. Here, pMNIST refers to the permuted MNIST variant (Goodfellow et al., 2013). It is clear from the results that using the proposed initializing technique provides positive improvements in test accuracy. Additionally, employing SII shows clear enhancement in training efficiency, achieving convergence at greater speed. This effect is depicted in Figure 6, where the test acccuracy curves are illustrated across the training epochs.
@@ 6.3 Multivariate Time-Series Forecasting
where $\hat{\bm{y}}\in\mathbb{R}^{{d_{y}}}$ is the measurement prediction, and $\bm{u}\in\mathbb{R}^{{d_{u}}}$ is a vector of input measurement signals. Importantly, the system is partly described by a state-transition function (19a), with unknown states and dynamics. In order to compute predictions according to (19b), the latent states $\bm{z}\in\mathbb{R}^{{d_{z}}}$ must be solved for, which requires that $\bm{z}_{0}$ is known. Inspired by Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014), we employ an encoder network $g_{\phi}$ that outputs the variational posterior over the initial state $\bm{z}_{0}\sim q_{\phi}(\bm{z}_{0}|\bm{u}_{0})$ and parametrize it as a diagonal Gaussian with learnable parameters.
@@ Results
A total of $3\times 10$ different models per initialization method were trained, consisting of $3$ random data splits (to mitigate some selection bias (Cawley & Talbot, 2010)) with $10$ different model configurations each. Test performance is presented in Table 3. Using SII results in a significant improvement in test performance compared to the default initialization. This enhancement is further supported by the faster convergence rate, as depicted in Figure 7.
@@ 8 Conclusions
This paper analyses how the stability properties of the system and stability regions of the chosen numerical solver affect the training and prediction performance of neural ODEs. Further, it is illustrated how standard techniques for the initialization of the network, without considering the stability properties of the system and solver, may lead to slow training and suboptimal performance. Based on this, a general stability-informed initialization (SII) method is developed that adapts to the stability region of a chosen solver. The effectiveness of the approach is demonstrated in various machine-learning tasks, including successful applications to real-world measurement data. In all cases, increased efficiency of training, less susceptibility to random fluctuations, and improved model performance are demonstrated.
@@ Acknowledgements
This research was supported by the Strategic Research Area at Linköping-Lund in Information Technology (ELLIIT) and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. The authors would like to thank the reviewers for their insightful comments and suggestions, which have significantly improved the manuscript.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.